write the function in terms of the cofunction of a complementary angle
tan5pi/12
thank you (:

Answer by Theo(13342)

90 degrees is equal to pi/2 which is equal to 6pi/12.
6pi/12 - 5pi/12 = 1pi/12.
that's the complementary angle to 5pi/12 because 5pi/12 + 1pi/12 = 6pi/12 which is equal to pi/2 which is equal to 90 degrees.
your co-function would be cot(pi/12).
tan(5pi/12) should be equal to cot(pi/12) if I'm correct.
use your calculator to get tan(5pi/12) and cot(pi/12)
cot(pi/12) is equal to 1/tan(pi/12).
tan(5pi/12) = 3.73205...
cot(pi/12) = 1/tan(pi/12) = 3.73205...
5/pi/12 * 180/pi = 5*180/12 = 5*15 = 75 degrees
set your calculator to degree mode.
tan(75) = 3.73205...
complementary angle of 75 degrees is 90 - 75 = 15 degrees.
cot(15) = 1 /tan(15) = 3.73205...
looks like we're good